OpenStudy (anonymous):

Question 4 Which of the following expressions is the conjugate of a complex number with −6 as the real part and −5i as the imaginary part? 6 + 5i 6 − 5i −6 − 5i −6 + 5i Question 5 Simplify 81 to the negative 3 fourths power −27 negative 1 over 27 27 1 over 27 Question 6 Find the solution of three times the square root of the quantity of x plus 5 equals negative 9 x = 4; not extraneous x = 4; extraneous x = 0; extraneous x = 0; not extraneous

OpenStudy (anonymous):

@phi

OpenStudy (math&ing001):

For the conjugate, you simply change the sign of the imaginary part : conj(-6-5i) = -6 +5i

OpenStudy (anonymous):

do you know 5 or 6?

OpenStudy (math&ing001):

\[81 = 3^{4} so (3^{4})^{-\frac{ 3 }{ 4 }}=3^{-3}=\frac{ 1 }{ 27 }\]

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

\[3\sqrt{x+5}=-9 so \sqrt{x+5}=(-3)^{2}=9 so x=4\]

OpenStudy (anonymous):

is it not extraneous or extraneous?

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

To know if it's extraneous or not, plug it in the first equation and see if it checks.

OpenStudy (anonymous):

no it doesn't

OpenStudy (anonymous):

so which one would it be?

OpenStudy (math&ing001):

So it's extraneous